# Asymptotic regularity of a generalised stochastic Halpern scheme

### v2

@@ Abstract.
We provide abstract, general and highly uniform rates of asymptotic regularity for a generalized stochastic Halpern-style iteration, which incorporates a second mapping in the style of a Krasnoselskii-Mann iteration. This iteration is general in two ways: First, it incorporates stochasticity in a completely abstract way rather than fixing a sampling method; secondly, it includes as special cases stochastic versions of various schemes from the optimization literature, including Halpern’s iteration as well as a Krasnoselskii-Mann iteration with Tikhonov regularization terms in the sense of Boţ, Csetnek and Meier. For these specific cases, we in particular obtain linear rates of asymptotic regularity, matching (or improving) the currently best known rates for these iterations in stochastic optimization, and quadratic rates of asymptotic regularity are obtained in the context of inner product spaces for the general iteration. At the end, we briefly sketch how the schemes presented here can be instantiated in the context of reinforcement learning to yield novel methods for Q-learning.
@@ 1. Introduction
Approximating fixed points of nonexpansive mappings is one of the most fundamental tasks in nonlinear analysis and optimization. The problem becomes particularly interesting when we only have noisy versions of those mappings, in which case the resulting approximation methods become stochastic processes. Concrete examples of this general situation include model-free reinforcement learning algorithms, where variants of $Q$-learning, for instance, can be viewed as stochastic methods for computing fixed points of nonexpansive operators.
To be more concrete, let $(X,\left\lVert\cdot\right\rVert)$ be a separable real-valued normed space and $T,U:X\to X$ be two nonexpansive mappings on $X$, i.e.
@@
for all $x,y\in X$. In order to approximate common fixed points of two such mappings under stochastic noise constraints, we introduce in this paper the so-called stochastic Halpern-Mann iteration, given by the schema
@@
Our main scheme (1) represents a stochastic analogue of the deterministic Halpern-Mann scheme for two operators, schematically appearing already in [37], and recently rediscovered in a much more generalized setting in [14]. This scheme integrates two of the most prominent methods for approximating fixed points for nonexpansive mappings, the Krasnoselskii-Mann method [24, 32] and Halpern’s method [16] (see also [42] and [43]), with the intended gain of combining the beneficial features of both, in particular the strong convergence of Halpern’s method even in infinite dimensional spaces. We conjecture that this combination gains further significance in the stochastic setting, where we anticipate that it can be used in particular to devise novel reinforcement learning algorithms or stochastic splitting methods.
@@
In contrast to the deterministic scheme, (1) is designed to capture situations in which one does not have direct access to $Tx_{n}$ (respectively $Uy_{n}$), but can only use noisy versions $\tilde{T}$ of $T$ (and $\tilde{U}$ of $U$). Intuitively, $\xi_{n}$ would then represent the difference between $Tx_{n}$ and the corresponding approximation of $Tx_{n}$ obtained from $\tilde{T}$ through a suitable sampling method (and similarly for $\delta_{n}$ and $Uy_{n}$), though our presentation is fully abstract and we will make no assumptions about $\xi_{n}$ or $\delta_{n}$, other than imposing controls on the means $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$, which however are carefully chosen so that they can be achieved in concrete scenarios through sampling methods such as minibatching under suitable assumptions. In particular, we make no independence assumptions on the noise terms $\xi_{n}$ and $\delta_{n}$.
@@
The main results of the paper establish general conditions under which we guarantee the asymptotic regularity of the scheme (1), both in the traditional sense of $\left\lVert x_{n}-x_{n+1}\right\rVert$ (sometimes called the discrete velocity [3]) and also relative to the mappings, i.e. considering the displacements $\left\lVert x_{n}-Tx_{n}\right\rVert$ and $\left\lVert x_{n}-Ux_{n}\right\rVert$. Furthermore, we establish these asymptotic regularity results both in expectation and almost surely, i.e. we show both
@@
$\left\lVert x_{n}-x_{n+1}\right\rVert,\ \left\lVert x_{n}-Tx_{n}\right\rVert,\ \left\lVert x_{n}-Ux_{n}\right\rVert\to 0\text{ almost surely}$
under suitable conditions on the auxiliary parameters as well as the noise. In the general case this requires a subtle geometric argument and involves a delicate interplay between the two modes of stochastic convergence, which manifest, respectively, in the additional assumptions that $X$ is uniformly convex and that the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is uniformly integrable. Most importantly, in all cases we provide explicit convergence rates for these expressions, under very general conditions and dependent only on a few moduli witnessing quantitative aspects of our main assumptions. In particular, we identify natural circumstances under which these rates are very fast, reaching up to linear speed in special cases.
To date, only the very simple instance of (1) corresponding to a stochastic variant of Halpern’s method [16] has been previously studied, where setting $U:=\mathrm{Id}$ and $\delta_{n}:=0$ gives rise to the iteration
$x_{n+1}=(1-\alpha_{n})(Tx_{n}+\xi_{n})+\alpha_{n}u,$
considered in [5] (for finite dimensional normed spaces), a scheme which in the Euclidean setting has recently received a great deal of attention in the context of stochastic monotone inclusion problems [6, 12, 31, 45]. In all cases, controlling the variance of the noise terms is crucial for convergence, and this is just one of several elements that makes the analysis of stochastic schemes such as (1) markedly different from that of their nonstochastic counterparts, some of the others being a focus on oracle complexity, and the relevance of stochastic methods to machine learning.
While we focus on the special case of the stochastic Halpern iteration (1) at several points in the paper, where it forms a useful example and where in particular our results on asymptotic regularity in that case generalize and improve those presented in [5], our method (1) is certainly not limited to this special case, and in line with [14] encompasses stochastic variants of other well-known deterministic methods.
An important example of this is represented by setting $T:=\mathrm{Id}$ and $\xi_{n}:=0$, as well as $u:=0$ and $\gamma_{n}:=1-\alpha_{n}$, whereby we obtain a version of the Krasnoselskii-Mann iteration with Tikhonov regularization terms considered in [2, 44] that now incorporates stochastic noise
$x_{n+1}=(1-\beta_{n})(U(\gamma_{n}x_{n})+\delta_{n})+\beta_{n}(\gamma_{n}x_{n})$
[…]
While the special cases of (1) and (1) exhibit linear rates of asymptotic regularity under very mild (quantitative) assumptions on the space, parameters and noise, the general iteration (1) proves to be more complex as mentioned before, depending to a large degree on the geometry of the space: In the general case, we assume $X$ to be uniformly convex, and the rate then depends on a corresponding modulus witnessing this uniform convexity. Our main geometric argument is carried out in a pointwise manner, yielding almost sure convergence (with rates). Nevertheless, the passage back to convergence in mean requires a uniform integrability assumption (which as we show can actually be imposed on the noise terms), and our rates then also depend on a corresponding modulus of uniform integrability, which we define and use for the first time in this paper. However, fast rates also apply to the general scheme (1), where in the special case of a uniformly convex space of power type $p\geq 2$, the complexity of those rates is of order $p$, and so in particular we get quadratic rates of asymptotic regularity for inner product spaces.
While all of our rates are novel, we emphasise that the “qualitative” asymptotic regularity results, namely convergence alone even without considering the quantitative aspects, are also to the best of our knowledge new, for any stochastic iteration contained in (1) that goes beyond the Halpern iteration.
@@
All of the results obtained in this paper are motivated via the methodology of the proof mining program, a subfield of mathematical logic which combines an abstract approach to proofs in mainstream mathematics with the extraction of computational information, such as bounds or rates, from those proofs. We refer to the seminal monograph [20] for a comprehensive overview of both theoretical as well as applied aspects of this program, along with the survey [22] for an overview of more recent applications to nonlinear analysis. Proof mining has been widely applied in nonlinear analysis, and has found particular success in providing quantitative convergence results for Halpern’s iteration and its many variants, with notable instances ranging from initial rates of asymptotic regularity for Halpern’s iteration given by Leuştean [27] and the first analysis of Wittmann’s proof of the strong convergence of Halpern’s iteration given in the seminal paper of Kohlenbach [21], to the extensions of these results to nonlinear context such as $\mathrm{CAT}(0)$-spaces as in [23] (by a logical analysis of a corresponding convergence proof by Saejung [39]). They also include extensions of the Halpern iteration [40] for the modified Mann iteration introduced in [18] (and extended to nonlinear spaces in [11]) as well as the Krasnoselskii-Mann iteration with Tikhonov regularization terms and its extensions as in [7, 8, 9, 13] (with [7] of particular note, as linear rates of asymptotic regularity are there obtained for the first time in the context of applications of proof mining). In particular, the definition of the deterministic Halpern-Mann method given in [14] and its corresponding convergence proof were motivated by these logical considerations, as were the recent rates of asymptotic regularity given for this iteration in [30].
@@
The present work departs from the aforementioned case studies in nonlinear analysis in that it incorporates, for the first time, stochasticity. In this way it forms part of a recent advance of proof mining into probability theory, which comprises both new developments in the logical foundations of probability theory due to first author and Neri [33], together with applied results on the quantitative aspects of stochastic processes by the authors and Neri [34, 35, 36]. In particular, the present paper is one of the first applications of proof mining to stochastic optimization: It represents a particularly interesting case study in this respect, in that it does not readily follow from analogous quantitative results in the deterministic setting (such as in [14, 30]), but requires a substantial arsenal of new quantitative ideas for this stochastic setting, as will be introduced and discussed in detail throughout the paper. We stress that while this logical perspective was crucial in obtaining the present results, the paper does not rely on any notions from logic at all.
However, the main motivation of this paper is not so much its novelty within the proof mining program, but the real applicability of the main results in concrete areas. For this reason, our last section is dedicated to a brief outlook on further results arising from the present work, in particular concerning oracle complexity and reinforcement learning, and in the context of the latter how present schemes can yield novel methods for Q-learning, all of which will be discussed in further detail a forthcoming paper. While not discussed there, we also want to mention that another promising application is represented by stochastic versions of the splitting methods with Tikhonov regularization terms as studied in [2], but we leave this for future work, as we do the study of the convergence of the method (1), where we expect that strong convergence results hold for this method also in infinite dimensional spaces, analogously to the deterministic case.

### v3

@@ Abstract.
We provide abstract, general and highly uniform rates of asymptotic regularity for a generalized stochastic Halpern-style iteration, which incorporates a second mapping in the style of a Krasnoselskii-Mann iteration. This iteration is general in two ways: First, it incorporates stochasticity completely abstractly, rather than fixing a sampling method; second, it includes as special cases stochastic versions of various schemes from the optimization literature, including Halpern’s iteration as well as a Krasnoselskii-Mann iteration with Tikhonov regularization terms in the sense of Boţ, Csetnek and Meier (where this stochastic variant of the latter is considered for the first time in this paper). For these specific cases, we obtain linear rates of asymptotic regularity, matching (or improving) the currently best known rates for these iterations in stochastic optimization, and quadratic rates of asymptotic regularity are obtained in the context of inner product spaces for the general iteration. We conclude by discussing how variance can be managed in practice through sampling methods in the style of minibatching, how our convergence rates can be adapted to provide oracle complexity bounds, and by sketching how the schemes presented here can be instantiated in the context of reinforcement learning to yield novel methods for Q-learning.
@@
Acknowledgments: We want to thank the two anonymous referees for their helpful suggestions which improved the paper in various places.
@@ 1. Introduction
1.1. Background and motivation
Approximating fixed points of nonexpansive mappings is one of the most fundamental tasks in nonlinear analysis and optimization. The problem becomes particularly interesting when we only have noisy versions of those mappings, in which case the resulting approximation methods become stochastic processes. Model-free reinforcement learning algorithms represent just one well-known example of this general situation, where variants of $Q$-learning, for instance, can be viewed as stochastic methods for computing fixed points of nonexpansive operators.
Let $(X,\left\lVert\cdot\right\rVert)$ be a separable real-valued normed space and $T,U:X\to X$ be two nonexpansive mappings on $X$, i.e.
@@ 1.1. Background and motivation
for all $x,y\in X$. In order to approximate common fixed points of two such mappings under stochastic noise constraints, we introduce in this paper the so-called stochastic Halpern-Mann iteration, given by the scheme
@@
Our main scheme (1.1) represents a stochastic analogue of the deterministic Halpern-Mann scheme for two operators (HM), schematically appearing already in [40], and recently (re-)discovered in a much more generalized setting in [13]. This scheme integrates two of the most prominent methods for approximating fixed points for nonexpansive mappings, the Krasnoselskii-Mann method [24, 35] and Halpern’s method [14] (see also [48] and [49]), with the intended gain of combining the beneficial features of both, in particular the strong convergence of Halpern’s method even in infinite dimensional spaces. We propose that this combination gains further significance in the stochastic setting, where we anticipate that it can be used in particular to devise novel reinforcement learning algorithms or stochastic splitting methods.
@@
In contrast to the deterministic scheme, (1.1) is designed to capture situations in which one does not have direct access to $Tx_{n}$ (respectively $Uy_{n}$), but can only use noisy versions $\tilde{T}$ of $T$ (and $\tilde{U}$ of $U$). Intuitively, $\xi_{n}$ would then represent the difference between $Tx_{n}$ and the corresponding approximation of $Tx_{n}$ obtained from $\tilde{T}$ through a suitable sampling method (and similarly for $\delta_{n}$ and $Uy_{n}$), though our presentation is fully abstract and we will make no assumptions about $\xi_{n}$ or $\delta_{n}$, other than imposing controls on $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$, in a way which can be easily achieved in concrete scenarios through sampling methods such as minibatching, as will also be discussed later.
To date, only the very simple instance of (1.1) corresponding to a stochastic variant of Halpern’s method (H) has been previously studied. Concretely, setting $U:=\mathrm{Id}$ and $\delta_{n}:=0$ gives rise to the iteration
$x_{n+1}=(1-\alpha_{n})(Tx_{n}+\xi_{n})+\alpha_{n}u,$
considered on an abstract level in [4] (for finite dimensional normed spaces), a scheme which in the Euclidean setting has recently received a great deal of attention in the context of stochastic monotone inclusion problems [5, 11, 31, 51]. In all cases, controlling the variance of the noise terms is crucial for convergence, and this is just one of several elements that makes the analysis of stochastic schemes such as (1.1) markedly different from that of their nonstochastic counterparts, some of the others being a focus on oracle complexity, and the relevance of stochastic methods to statistics and machine learning.
While we focus on the special case of the stochastic Halpern iteration (1.1) at several points in the paper, where it forms a useful example, our method (1.1) is certainly not limited to this special case, and in line with [13] encompasses stochastic variants of other well-known deterministic methods. An important example of this is represented by setting $T:=\mathrm{Id}$ and $\xi_{n}:=0$, as well as $u:=0$ and $\gamma_{n}:=1-\alpha_{n}$, whereby we obtain a version of the Krasnoselskii-Mann iteration with Tikhonov regularization terms (KM-T) considered in [1, 50] that now incorporates stochastic noise, taking the form of
$x_{n+1}=(1-\beta_{n})(U(\gamma_{n}x_{n})+\delta_{n})+\beta_{n}(\gamma_{n}x_{n}).$
Alternatively, the method can be seen as the stochastic Krasnoselskii-Mann iteration as considered in [3] with Tikhonov regularization terms as considered in [1]. This method is known to produce fast asymptotic behavior in the deterministic setting, and in contrast to the standard Krasnoselskii-Mann scheme, also benefits from strong convergence results similarly to Halpern’s iteration [1]. As we will show in this paper, in particular the first of these features extends to the stochastic setting (while we also lay the foundation for the second, as discussed later). For a simple overview, we present the relationships between the different stochastic schemes considered here, as well as their deterministic counterparts, in Figure 1.
[…]
Figure 1. Relationships of the stochastic Halpern-type schemes and their deterministic variants
To our knowledge, each of the stochastic schemes (1.1) and (1.1) are introduced here for the first time.
1.2. Overview of the main results
Our main results comprise both abstract convergence theorems for the stochastic schemes in Figure 1, valid under very broad assumptions (and with convergence rates given at a corresponding level of generality), along with a series of special cases where fast, linear rates are possible, the latter demonstrating that our framework includes stochastic algorithms that come with state-of-the-art convergence guarantees of a kind only recently established for deterministic Halpern-type schemes [32, 43]. We now outline each of these in turn.
1.2.1. General (quantitative) asymptotic regularity results
@@
The main results of the paper establish general conditions under which one can guarantee the asymptotic regularity of the scheme (1.1), both in the traditional sense of $\left\lVert x_{n}-x_{n+1}\right\rVert$ (sometimes called the discrete velocity [2]) and also relative to the mappings, i.e. considering the displacements $\left\lVert x_{n}-Tx_{n}\right\rVert$ and $\left\lVert x_{n}-Ux_{n}\right\rVert$. Furthermore, we establish these asymptotic regularity results both in expectation and almost surely, that is we show both
@@
$\left\lVert x_{n}-x_{n+1}\right\rVert,\ \left\lVert x_{n}-Tx_{n}\right\rVert,\ \left\lVert x_{n}-Ux_{n}\right\rVert\to 0\text{ almost surely}.$
Most importantly, in all cases we provide explicit convergence rates for these expressions. In the case of convergence in mean, these rates take the form of functions $\varphi$ which guarantee
@@
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]<\varepsilon\text{ for all }\varepsilon>0\text{ and any }n\geq\varphi(\varepsilon),$
and similarly for $\mathbb{E}[\left\lVert x_{n}-Tx_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert x_{n}-Ux_{n}\right\rVert]$. In the case of our almost-sure convergence results, our rates instead take the form of functions $\Phi$ which guarantee
$\mathbb{P}(\exists n\geq\Phi(\lambda,\varepsilon)(\left\lVert x_{n}-x_{n+1}\right\rVert\geq\varepsilon))<\lambda$
for all $\varepsilon,\lambda>0$ (and similarly for $\left\lVert x_{n}-Tx_{n}\right\rVert$ and $\left\lVert x_{n}-Ux_{n}\right\rVert$). In all cases, these rates are explicitly constructed under very general conditions and depend only on a few moduli witnessing quantitative aspects of our main assumptions.
These general asymptotic regularity results and their corresponding rates for the scheme (1.1), both in the sense of the discrete velocity and relative to the mappings, can be found in Theorems 3.23.5. While the present paper features various fine-grained discussions on the precise assumptions necessary for each of these results, they essentially amount to
  • the existence of a common fixed point of $T$ and $U$,
  • standard conditions on the parameters $(\alpha_{n})$ and $(\beta_{n})$, and
  • variance control in the form of $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<+\infty$ and $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<+\infty$.
In their quantitative variants, the respective rates correspondingly rely on bounds and rates witnessing these assumptions in various ways. While the results on the discrete velocity as given in Theorems 3.2 and 3.3 are proved unconditionally, asymptotic regularity relative to the mappings turns out to be more involved, and the respective Theorems 3.4 and 3.5 depend a priori on the assumption that $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$, either in mean or almost surely. It is precisely at this point that our treatment divides according to the main schemes we consider in Figure 1. For (1.1) and (1.1), this premise either trivializes or can be directly derived without additional conditions (see Theorem 3.8), but for the general case of (1.1), establishing $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely requires a subtle pointwise geometric argument based on the additional assumption that the underlying space $X$ is uniformly convex (see Theorem 3.12). Our corresponding result in mean (see Theorem 3.18) is effectively obtained as a direct lift of the pointwise argument, which we achieve in an abstract way by showing that the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is uniformly integrable given the existence of a common fixed point and the variance assumptions $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<+\infty$ and $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<+\infty$.
Naturally, these geometric and probabilistic considerations also influence the corresponding rates, where the crucial assumption of uniform convexity and passage through uniform integrability result in a dependence of the rate on a modulus measuring the degree of uniform convexity of $X$, along with a similar modulus measuring the degree of uniform integrability of $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. In particular, we show how such moduli can be explicitly constructed for particular spaces, and from natural quantitative integrability assumptions on the error terms, respectively.
To the best of our knowledge, all the respective rates for (1.1) and (1.1), as well as even the “qualitative” asymptotic regularity results, that is convergence alone even without considering the quantitative aspects, are completely novel. In terms of complexity and structure, they seem to match previously constructed general rates for their deterministic analogs (KM-T) and (HM) derived in [6, 7, 8, 12] and [13, 30], respectively, with the additional component of a modulus of uniform integrability featuring in the case of (1.1) as mentioned before. In the case of the stochastic Halpern iteration (1.1), the only scheme already to have been studied, our convergence theorems are the best yet in that they both generalize and improve the asymptotic regularity results recently presented in [4].
1.2.2. Fast rates of asymptotic regularity
Beyond just presenting general and rather abstract convergence results and rates as outlined above, we also identify natural circumstances under which these rates are very fast, reaching up to linear speed in special cases. For the parameter choices $\alpha_{n}=2/(n+2)$ and $\beta_{n}=\beta\in(0,1)$, together with sufficiently fast decaying errors with $\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K/(n+2)^{2}$ for some constant $K$, we derive (see Theorem 4.4) linear rates for the discrete velocity of the general scheme (1.1) in the form of
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]\leq\frac{K^{\prime}}{n+2}\text{ and }\mathbb{P}\left(\exists i\geq n\left(\left\lVert x_{i}-x_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{K^{\prime}}{n+2}$
for all $n\in\mathbb{N}$ and $\varepsilon>0$, and some suitable (explicitly definable) constant $K^{\prime}$. Similar results also hold for the discrete velocity of $(y_{n})$ in mean and almost surely (see Theorem 4.6). In particular, we want to note that the above results, as well as all other linear rates, are actually given in the form of nonasymptotic guarantees, also in the case of the almost sure rates.
In the case of asymptotic regularity relative to the mappings, our results again become sensitive to the choice of the scheme: The special cases of (1.1) and (1.1) exhibit linear nonasymptotic guarantees under the same assumptions on the parameters as detailed above. In the case of (1.1) (see Theorem 4.8), we in particular have
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\leq\frac{K^{\prime}}{n+2}\text{ and }\mathbb{P}\left(\exists i\geq n\left(\left\lVert Tx_{i}-x_{i}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{K^{\prime}}{n+2}$
for all $n\in\mathbb{N}$ and $\varepsilon>0$ and some suitable (explicitly definable) constant $K^{\prime}$ similar to before. Effectively the same results hold for $U$ in place of $T$ in the case of (1.1) (see Theorem 4.9).
The results for the general scheme (1.1) again rely both on assumptions on the geometry of the space $X$ and the uniform integrability of the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. In the special case of a uniformly convex space of power type $p\geq 2$, our results in particular yield rates of order $p$ so that we obtain quadratic rates in the case of inner product spaces, with order
$\mathcal{O}(\varepsilon^{-2}\lambda^{-2})\text{ or }\mathcal{O}(\varepsilon^{-2}\mu(\varepsilon/4)^{-2})$
for $\left\lVert Ux_{n}-x_{n}\right\rVert,\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely or in mean, respectively, where $\mu$ is our modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ (see Theorem 4.12 and Remark 4.13).
As with our generally constructed rates, our fast rates for the schemes (1.1) and (1.1) are to the best of our knowledge also novel. In regards to complexity, and seem to match the complexities observed in the deterministic case (we again refer to [13, 30] and [6, 7, 8, 12], respectively). For the scheme (1.1), our results improve the fast rates presented in [4] by removing the respective logarithmic factors contained therein. In particular, by restricting our attention to the deterministic Halpern iteration (H), we reobtain the linear rates of asymptotic regularity previously derived in [32, 43] (albeit with different constants), which are known to be tight (see [32]). Indeed, our linear rates are derived by adapting the approach of [43]). Based on their inherent similarity, it can moreover be believed that the linear rates of asymptotic regularity observed for the deterministic Krasnoselskii-Mann iteration with Tikhonov regularization terms (KM-T) are also tight, in which case our corresponding results for the stochastic scheme (1.1) would be tight as well, although we are not aware of any results that prove this explicitly. The question for tight rates for the general Halpern-Mann scheme, both in its deterministic variant (HM) as well as in the stochastic version (1.1) presented here, remains an open problem.
1.3. Future applications and developments
Beyond our series of general theoretical results, one of the core motivations for this paper is the real applicability of those results in concrete areas, and we conclude with a section on how this might be achieved. First, we discuss how the requisite variance control can be managed in a practical way through the use of sampling techniques such as minibatching, and connected with this we show how our complexity results can be lifted to corresponding results on oracle complexity. We then outline several concrete applications of our methods, sketching the particularly interesting case of reinforcement learning, where the novel schemes (1.1) and (1.1) can be instantiated in the style of Q-learning, as done recently for the stochastic Halpern iteration in [4]. A proper account of these applications will be provided in a forthcoming paper by the authors. Another crucial property of Halpern-type iterations not addressed in the present paper is the fact that they remain strongly convergent also in infinite dimensional spaces. These strong convergence results extend to the stochastic setting, and will be similarly addressed in forthcoming work which in particular relies on the asymptotic regularity results established here.
1.4. This paper in connection with the proof mining program
@@
All of the results obtained in this paper are motivated via the methodology of the proof mining program, a subfield of mathematical logic which combines an abstract approach to proofs in mainstream mathematics with the extraction of computational information, such as bounds or rates, from those proofs. We refer to the seminal monograph [19] for a comprehensive overview of both theoretical as well as applied aspects of this program, along with the survey [21] for an overview of more recent applications to nonlinear analysis. Proof mining has been widely applied in nonlinear analysis, and has found particular success in providing quantitative convergence results for Halpern’s iteration and its many variants, with notable instances ranging from initial rates of asymptotic regularity for Halpern’s iteration given by Leuştean [27] and the first analysis of Wittmann’s proof of the strong convergence of Halpern’s iteration given by Kohlenbach [20], to the extensions of these results to nonlinear context such as $\mathrm{CAT}(0)$-spaces as in [22] (by a logical analysis of a corresponding convergence proof by Saejung [44]). They also include extensions of the Halpern iteration [45] for the modified Mann iteration introduced in [16] (and extended to nonlinear spaces in [10]) as well as the Krasnoselskii-Mann iteration with Tikhonov regularization terms and its extensions as in [6, 7, 8, 12] (with [6] of particular note, as linear rates of asymptotic regularity are there obtained for the first time in the context of applications of proof mining). In particular, the definition of the deterministic Halpern-Mann method given in [13] and its corresponding convergence proof were motivated by these logical considerations, as were the recent rates of asymptotic regularity given for this iteration in [30].
@@
The present work departs from the aforementioned case studies in nonlinear analysis in that it incorporates, for the first time, stochasticity. In this way it forms part of a recent advance of proof mining into probability theory, which comprises both new developments in the logical foundations of probability theory due to first author and Neri [36], together with applied results on the quantitative aspects of stochastic processes by the authors and Neri [37, 38, 39]. In particular, the present paper is one of the first applications of proof mining to stochastic optimization, and the very first to consider a concrete stochastic algorithm. It represents a particularly interesting case study in this respect, in that it does not readily follow from analogous quantitative results in the deterministic setting (such as in [13, 30]), but requires a substantial arsenal of new quantitative ideas for this stochastic setting. These include quantitative, stochastic variants of a crucial abstract lemma on recurrence inequalities due to Xu [49] (Lemma 2.4), including an adaptation of the “fast” variant due to Sabach and Shtern [43] (Lemma 4.2). Furthermore, proof mining also motivated the definition of a modulus of uniform integrability and its use as a suitable assumption to derive rates of asymptotic regularity in expectation in the general case of the iteration (1.1). We envisage that all of these tools will be relevant in subsequent applications of proof mining in stochastic optimization, just as their deterministic counterparts have been used repeatedly for proof mining in nonstochastic optimization.
We stress that while this logical perspective was crucial in obtaining the present results, the paper does not rely on any notions from logic at all.